OpenStudy (anonymous):

Solve the equation. 3(4x - 3) - 19 = 8x - 4

OpenStudy (kira_yamato):

12x - 9 = 8x + 23 4x = 32 x = 8

OpenStudy (pradipgr817):

12x-28=8x-4 4x=24 x=6

OpenStudy (anonymous):

3(4x-3) - 19 = 8x -4 12x - 9 - 19 = 8x - 4 12x - 28 = 8x - 4 12x = 8x + 24 4x = 24 x = 6

OpenStudy (anonymous):

3(4x - 3) - 19 = 8x - 4 => 12x-9-19 =8x -4 => 12x-8x = 28-4 => 4x=24 => x=4